# HeadEvolver: Text to Head Avatars via Expressive and Attribute-Preserving Mesh Deformation

### v2

@@ 4 Experiments
In this section, we present experiments to evaluate the efficacy of our method both quantitatively and qualitatively for text-to-avatar creation. Then, we demonstrate the results of the ablation study that validate the significance of our key designs for vector field and regularization terms in Sec. 4.3.
Implementation details. For geometry and texture generation, we render normal and shaded images at a resolution of $512\times 512$ pixels and feed them into the Realistic Vision 5.1 (RV 5.1) [24]. Compared with Stable Diffusion 2.1 [38], we find that RV 5.1 supports a more appealing avatar appearance. The multi-face Janus and texture misalignment problems are further alleviated by introducing ControlNetMediaPipeFace [57]. The generation process runs on a single NVIDIA RTX $4090$ GPU with $20,000$ iterations per avatar, which takes around $70$ minutes. ResAdapter [7] is further utilized to generate the albedo map with a 1024 $\times$ 1024 resolution. Negative prompts are directly leveraged to enhance the realism and details of the texture, e.g., replacing the empty prompt in classifier-free guidance (CFG) with “unrealistic,” “blurry,” “low quality,” “saturated,” “low contrast,” etc. In addition to the proposed regularization terms, we enforce symmetry along the y-axis in each iteration to constrain the overall position offsets of the generated mesh.
@@ 4.1 Qualitative Evaluation
Our evaluation compares two explicit mesh manipulation methods from TADA [29] and TextDeformer [10], and one implicit method from Fantasia3D [6] as baselines. We do not compare with HeadArtist [31] and DreamFace [56] due to their lack of open-sourcing and primary focus on rendering quality. Instead, our framework emphasizes on generating high-quality mesh topologies. We utilized FLAME’s neutral expression as a unified input template mesh to conduct experiments. Visual comparison results are in Fig. 3. TADA and Fantasia3D demonstrate their capabilities of detail-preserving geometry with 3D head priors. However, their results still suffer from geometric artifacts in semantically dense areas, such as noisy vertices around eye regions, although this is less noticeable in a textured rendering. Our method mitigates this issue with the spatial alignments guided by landmarks and opacity regularization terms. Compared to TextDeformer, our approach exhibits similar mesh quality and amplifies facial characteristics following text specifications via proposed vector fields.
In summary, our framework generates expressive human head meshes that inherit facial attributes. The texture generation aligns with the geometry to create 3D assets ready for graphics applications. As depicted in Fig. 8 and Fig. 9, with text guidance, generated models from our method could be effectively manipulated to change appearance. Fig. LABEL:fig:teaser(e) and Fig. 14 illustrate the potential for seamless utilization of these avatars in 3D software by artists, as they exhibit distinct facial features and fewer artifacts.
Table 2: Quantitative comparison of state-of-the-art methods. The geometry CLIP score is measured on the shading images rendered with uniform albedo colors, the appearance CLIP score is evaluated on the images rendered with textures, and the self-intersection is quantified as a ratio between the number of self-intersections in the deformed mesh and the number of faces. $\uparrow$ indicates higher is better and $\downarrow$ indicates lower is better.
Metric Ours TADA TextDeformer Fantasia3D
Geometry CLIP $\uparrow$ 0.2506 $0.2475$ $0.2462$ $0.1815$
Appearance CLIP $\uparrow$ 0.2955 $0.2931$ $0.2718$
Self-Intersection $\downarrow$ 2.08% $14.22\%$ $2.19\%$
@@ 4.2 Quantitative Evaluation
We quantitatively evaluated our framework in generation consistency with text input and generation quality.
Generation Consistency with Text. We first evaluated the relevance of generated results to the corresponding text descriptions by calculating the CLIP score [37]. In addition, to separate the geometric attributes from the influence of texture, we rendered the generated geometry with a uniform albedo. Subsequently, we calculated the geometry CLIP score as a supplementary measure for evaluating the generation consistency of geometry with text prompts. We rendered 20 distinct views and computed the average scores respectively, as shown in Table 2. Our approach achieves the highest scores compared to the baseline methods.
Mesh Quality. We calculated the metric of self-intersection to measure mesh quality, which is the ratio of the number of self-intersections over the number of faces in the deformed mesh. Since the mesh from Fantasia3D is extracted through the marching cubes algorithm [32], it may produce skinny triangles containing small edges and angles with no self-intersections. Therefore, we only compared our method with TADA and TextDeformer. Unlike our method, geometry manipulation through direct vertex displacement tends to converge to a local minimum and disregards the triangulation of the shape.
User Study. We conducted a user study to evaluate the robustness and expressiveness of our method with 10 distinct text prompts, 4 of which are from TADA. We used a Google Form to assess 1) geometry quality, 2) texture quality, and 3) consistency with text. We recruited 46 participants, of whom 26 are graduate students majoring in computer graphics and vision, and 20 are company employees specializing in AI content generation. In this form, the participants were instructed to choose the preferred renderings of head avatars from different methods in randomized order, as shown in Table 3. The results show that participants preferred our method by a significant margin compared with the baseline methods in all three metrics.
Table 3: User evaluation of generated head avatars. Compared to other methods, ours is preferred by over 65% users in terms of geometry quality, texture quality, and consistency with text.
Preference $\uparrow$ Ours TADA TextDeformer Fantasia3D
Geometry Quality 65.9% $3.9\%$ $29.8\%$ $0.4\%$
Texture Quality 65.2% $31.9\%$ $2.9\%$
Consistency with Text 66.9% $28.9\%$ $2.3\%$ $1.9\%$
@@ 4.3 Ablation Study
Effect of Vector field. We conducted three sets of qualitative comparisons to demonstrate the effectiveness of the per-triangle vector field $H$. As shown in Fig. 5, using Jacobian $J$ alone tends to overly smooth the gradients, leading to a loss of distinctive character features. Although increasing the learning rate of $J$ highlights character identities, generated meshes are distorted and suffer from severe self-intersections. Furthermore, our experiments introduce the scalar field $S$ which represents the isotropic scaling by setting diagonal elements of each $H_{i}$ to equal values. The results indicate that anisotropic scaling guided through $H$ generates avatars bearing more geometry details and character features than those through $S$ (e.g., the eyebrow of ”Octopus Alien” and the ear of ”Goblin”). Furthermore, we demonstrate the tuning process of the learning rate (LR) to evaluate that the addition of $H$ effectively enhances the stability and expressiveness of deformations when the LR of $J$ is set in the reasonable range (e.g., between $0.01$ to $0.025$). As shown in Fig. 4, we provide the best LR settings, which are $0.015$ for $H$ and 0.01 for $J$ respectively. We also demonstrate the generality of vector field by applying it to the object generation framework (Fig. 16) and observe that the use of $H$ can make optimization converge to desired deformation states faster than using Jacobians only.
Regularization. In deformation optimization, facial features and head poses may deviate unexpectedly. Landmark regularization constrains generated meshes to the input shape’s canonical space, and contour-based regularization contributes to reasonable head shapes. We demonstrate the effectiveness of our regularization designs with another ablation study, as shown in Fig. 6.
Figure 4: Learning rate (LR) tuning. Supported by our experiments, adding per-triangle $H$ through $W$ can enhance the identity and character features of the avatar from the text prompt. We also observe the best LR of $W$ could be $0.015$, and larger LRs can cause unexpected severe deformations.
Figure 5: Ablation study on Vector Fields. The per-triangle $H$ can enhance the identity and character features of the avatar.
Figure 6: Ablation study results on regularization. Introducing an opacity map and landmark regularization contributes to maintaining symmetric facial features and aligning them with the pose of the input mesh.
@@ 5 Applications
Our method can produce head avatars that can be smoothly manipulated (e.g., animated via 3DMMs and text-based editing) and are compatible with graphics tools such as Blender and Maya. Compared with implicit representations, our method enables artists to interactively create, edit, and animate digital head avatar assets for various downstream applications.
Local geometry editing. Besides manipulating generated meshes in 3D software as artists usually do, our approach also supports efficient head editing with text descriptions (Figs. LABEL:fig:teaser(b) and  8). Only $H_{i}$ on certain parts is optimized given the text guidance.
Texture editing through text guidance. Our approach encourages independent texture generation on a fixed geometry, which enables texture editing specified by the text input (Fig. 9). Some unexpected color artifacts may be introduced due to the noises from the stable diffusion model.
Avatar animation. As shown in Figs. LABEL:fig:teaser(c) and  10, we demonstrate that the rigging and blendshapes of the head mesh model after deformation are well-preserved for creating head animations and shape manipulations.
Avatar generation from other template meshes. As described in Sec. 3, our approach is not constrained by specific mesh models (e.g., FLAME) as long as they are manifold mesh. Fig. 15 shows an example of generating a “Superman” head from crafted 3D assets even without labeling 3D facial landmarks.
@@ 6 Limitations and Future Work
Our framework has a few limitations. Although many parametric head models do not contain eyeball mesh intrinsically (e.g., Facescape [48], NPHM [12], and BFM [11]), our experiments removed the eyeball mesh from FLAME and SMPL-X in preprocessing since Jacobians rely on the manifold mesh structure and continuous mapping. A cage-based representation could offer the possibilities of deformations among multiple mesh instances [49], allowing for separately editable details like hair and accessories. Another minor limitation is that although we jointly optimize the geometry and texture, lighting is not disentangled from diffuse colors, which may add pixel-level noises and oversaturation artifacts on the albedo maps.
As Fig. 3 and Table 2 show, although our results achieve relatively good quantitative and qualitative results, our method cannot always generate corresponding geometry and texture. For example, eyeball appearance is produced only from the texture. Another future direction is to learn 3D head generation from a spherical mesh without strong avatar priors [34]. Designing an encoder to guide deformations with topology awareness may be necessary.
@@ 7 Conclusion
In this paper, we have introduced a text-guided generative framework for creating stylized 3D head avatars through learnable deformations. We use Jacobians as an intermediate representation for vertex displacements and improve the expressiveness of deformations by equipping Jacobians with learnable vector fields. As a result, such learnable deformations are embedded into score distillation for geometry and appearance generations. Besides deformation enhancement, proposed regularization terms help retain 3D attributes from the source mesh to support smooth reusability and editing by artists. Experiments on diverse text prompts demonstrate the significant performance of our method. Our method can potentially inform future research that better integrates neural generative techniques with practical graphics pipelines.

### v3

@@ 4 Experiments
In this section, we conduct experiments to evaluate the efficacy of our method both quantitatively and qualitatively for text-to-avatar creation. We then present two ablation studies that validate the significance of our key designs for vector fields and regularization terms in Section 4.4.
4.1 Implementation Details
For geometry and texture generation, we render normal and shaded images at a resolution of $1024\times 1024$ pixels and feed them into the Realistic Vision 5.1 (RV 5.1) [25]. Compared with Stable Diffusion 2.1 [42], we find that RV 5.1 supports a more appealing avatar appearance. The multi-face Janus and texture misalignment problems are further alleviated by introducing ControlNetMediaPipeFace [63]. The generation process runs on a single NVIDIA RTX $4090$ GPU with $20,000$ iterations per avatar, which takes around $30$ minutes. Negative prompts are leveraged to enhance the realism and details of the texture, e.g., “unrealistic,” “blurry,” “low quality,” “saturated,” “low contrast,” etc. Along with the proposed regularization terms, we enforce symmetry along the y-axis in each iteration to ensure a reasonable face appearance. Following established practices in TADA and DreamFace, the eyeball mesh is removed and can be manually reintegrated after the generation.
Figure 3: Qualitative appearance comparisons. Our method can produce diverse textured avatars.
Figure 4: Qualitative geometry comparisons. Our method effectively preserves the topology and semantics of the input template mesh, resulting in high-quality mesh models for smooth manipulations.
@@ 4.2 Qualitative Evaluation
Our evaluation compares two explicit mesh manipulation methods (i.e., TADA [30] and TextDeformer [9]), and two implicit methods (i.e., Fantasia3D [6] and HumanNorm [20]) as baselines. While our framework focuses on high-quality mesh topologies, we also compare rendering appearance with recent non-mesh-dependent methods [32, 14, 65] in supplemental material (Figure 19). We utilized FLAME [28] as a unified input template mesh to conduct the experiments. We present the visual comparison results in Figures 3 and 4. TADA and Fantasia3D show detail-preserving geometry with 3D head priors but still undergo geometric artifacts in dense areas like eye regions, which are less noticeable in textured rendering. HumanNorm performs underperforms in humanoid cases, e.g., “Terracotta Army” and “Monkey.” Compared to TextDeformer, our approach exhibits similar mesh quality and amplifies facial characteristics following text specifications.
In summary, our framework generates expressive head meshes that inherit facial attributes. The texture generation maintains decent geometry consistency, facilitating usage in graphics applications (Figures LABEL:fig:teaser(c) and 3).
@@ 4.3 Quantitative Evaluation
We quantitatively evaluated our framework for generation consistency with text input and generation quality.
Generation Consistency with Text. We first evaluated the relevance of generated results to the corresponding text descriptions by calculating the CLIP score [41]. We rendered 20 distinct views and computed the average scores respectively, as shown in Table 2. Our approach achieves the highest scores compared to the baseline methods.
Table 2: Quantitative comparison of state-of-the-art methods. The geometry CLIP score is measured on the shaded images rendered with uniform albedo colors [40], the appearance CLIP score is evaluated on the images rendered with textures, and the self-intersection is quantified as the ratio of self-intersected mesh faces to the total number of mesh faces.
Geometry Appearance Mesh
CLIP Score $\uparrow$ CLIP Score $\uparrow$ Self-Intersection $\downarrow$
Fantasia3D $0.1815$ $0.2718$
HumanNorm $0.2443$ $0.3031$
TextDeformer $0.2462$ $2.19\%$
TADA $0.2475$ $0.2931$ $14.22\%$
Ours 0.2538 0.3089 2.08%
Table 3: User evaluation of generated head avatars.
User Geometry Texture Consistency
Preference $\uparrow$ Quality Quality with Text
Fantasia3D $0.4\%$ $2.0\%$ $1.9\%$
HumanNorm $5.6\%$ $22.7\%$ $6.7\%$
TextDeformer $24.2\%$ $2.3\%$
TADA $3.9\%$ $13.1\%$ $22.2\%$
Ours 65.9% 62.2% 66.9%
Mesh Quality. We evaluate mesh quality through self-intersection. Since the mesh from Fantasia3D and HumanNorm are extracted through the marching cubes algorithm [33], the produced triangles have no self-intersections. Therefore, we only compared our method with TADA and TextDeformer. Unlike our method, direct vertex displacement tends to converge to a local minimum and disregards the triangulation of the mesh.
Figure 5: Ablation study on the vector fields. The per-triangle $H$ can enhance the identity and character features of the avatar.
Figure 6: Learning rate (LR) tuning. Supported by our experiments, adding per-triangle $H$ through $W$ can enhance the identity and character features of the avatar from the text prompt.
Figure 7: Preserved semantic consistency (e.g., 3D landmarks and dense segmentation) with our deformation-based framework.
User Study. We conducted a user study to evaluate the robustness and expressiveness of our method with 18 distinct text prompts, 6 of which are from TADA. We used a Google Form to assess 1) geometry quality, 2) texture quality, and 3) consistency with text. We recruited 46 participants, of whom 26 are graduate students majoring in computer graphics and vision, and 20 are company employees specializing in AI content generation. In this form, the participants were instructed to choose the preferred renderings of head avatars from different methods in randomized order, as shown in Table 3. The results show that participants preferred our method by a significant margin (over $60\%$).
Figure 8: Avatar generation with various template mesh inputs. The meshes from (a) an arbitrary head model and (b) SMPL-X [36] are deformed to align text descriptions respectively.
Figure 9: Generalization experiments. We implement vector fields in TextDeformer [9] and find that our method accelerates mesh optimizations to approach text guidance in early epochs.
@@ 4.4 Ablation Study
Effect of Vector field. We conducted three sets of qualitative comparisons to demonstrate the effectiveness of the per-triangle vector field $H$. As shown in Figure 5, using Jacobian $J$ alone tends to overly smooth the gradients, leading to a loss of distinctive character features. Increasing $J$’s learning rate emphasizes character identities but distorts meshes and causes self-intersections. Additionally, our experiments introduce the scalar field $S$, which represents the isotropic scaling by setting diagonal elements of each $H_{i}$ to equal values. The results indicate that anisotropic scaling guided through $H$ generates avatars bearing more geometry details and character features than those through $S$ (e.g., the eyebrow of ”Octopus Alien” and the ear of ”Goblin”). Furthermore, we demonstrate the tuning process of the LR to evaluate that the addition of $H$ effectively enhances the stability and expressiveness of deformations when the LR of $J$ is set in the reasonable range (e.g., between 0.01 to 0.025). In Figure 7, we provide the best LR settings, which are $0.015$ for $H$ and 0.01 for $J$. We also demonstrate the versatility of the vector field by applying it to the object generation framework (Figure 9), observing that the use of $H$ accelerates optimization convergence to desired deformation states compared to using Jacobians alone.
Regularization. In deformation optimization, facial features and head poses may deviate unexpectedly. Landmark regularization constrains generated meshes to the input shape’s canonical space, and contour-based regularization contributes to reasonable head shapes. We demonstrate the effectiveness of our regularization designs in Figure 11.
Figure 10: Ablation study on regularization. The contour-based and landmark regularizations contribute to maintaining symmetric facial features and alignment with the pose of the input mesh.
Figure 11: Our deformation-based framework offers effective geometry consistency for further 3D applications.
@@ 5 Applications
Our method produces head avatars that can be smoothly manipulated (e.g., animated via 3DMMs and text-based editing) and are compatible with graphics tools such as Blender and Maya. Compared with implicit representations [6, 14], our method enables artists to animate and edit head avatar assets for downstream applications (Figures LABEL:fig:teaser(c) and 16).
Local geometry editing. Besides manipulating generated meshes in 3D software as artists usually do, our approach also supports efficient head editing with text descriptions (Figures LABEL:fig:teaser(b) and 16). Only $H_{i}$ on certain parts is optimized given the text guidance.
Texture editing through text guidance. Our approach encourages independent texture generation on a fixed geometry, which enables texture editing specified by the text input (Figure 19). Some unexpected color artifacts may be introduced due to the noises from the diffusion model.
Avatar animation. As shown in Figures LABEL:fig:teaser(c) and 12, we demonstrate that the rigging and blendshapes of the head mesh model after deformation are well-preserved for creating head animations and shape manipulations.
Avatar generation from other template meshes. As detailed in Section 3, our method accommodates manifold meshes, not just specific models like FLAME. Figure 9 exemplifies a ”Superman” head creation without labeled 3D facial landmarks, using crafted 3D assets.
@@ 6 Limitations and Future Work
Our framework has a few limitations. Although many parametric head models do not contain eyeball mesh intrinsically (e.g., Facescape [52], NPHM [11], and BFM [10]), our experiments removed the eyeball mesh from FLAME and SMPL-X since Jacobians rely on the manifold mesh structure [39] (Figure 4.1). A cage-based representation may offer deformations among multiple mesh instances simultaneously [53], enabling separately editable details like hair and accessories. Another minor limitation exists in that lighting isn’t disentangled from diffuse colors, potentially causing noise and oversaturation artifacts.
It is worth mentioning that the generated avatars exhibit relatively good geometry consistency with texture (Figure 11), but our method cannot guarantee an exact match between geometry and texture. For instance, eyeball appearance only shows in the texture, a problem that also occurs in TADA and HumanNorm. Stronger semantic labeling (e.g., canny edges [37]) may mitigate this issue. Another future focus is to learn head generation from a 3D sphere without any avatar shape priors [35].
Figure 12: Example of attribute inheritance from FLAME [28]. If the input template mesh is from 3DMMs [4], users could manipulate the deformed mesh freely via morphable attributes such as expression and shape parameters.
@@ 7 Conclusion
In this paper, we have presented a text-guided generative framework for crafting stylized 3D head avatars through learnable deformations. By employing Jacobians as an intermediate representation for vertex displacements and enhancing them with a learnable vector field, our method has improved the expressiveness of the deformed mesh. Consequently, learnable deformations are embedded into score distillation for geometry and appearance generation. Our regularization terms help retain 3D attributes from the source mesh to support seamless editing by artists. Experiments with various text prompts exhibit the remarkable performance of our framework. Our work can inform future research on effectively integrating mesh-based neural techniques into practical graphics pipelines, such as boosting the 2D and 3D deformation ability by applying our method to APAP [55].
@@ 8 Acknowledgement
Supplemental
The supplementary material includes additional generation results and showcases applications such as text-guided geometry and texture editing.
Figure 13: Examples of local geometry editing. Our framework uses per-vertex segmentation information (Figure 7), enabling text-based mesh editing to alter avatars’ local features.
Figure 14: More representative generation results.
Figure 15: A limitation of our method. For exaggerated figures like “Hulk”, (a) though it does not impact further applications too much shown in Figure LABEL:fig:teaser(c), perfect geometry-texture alignment is not guaranteed. (b) By fixing the geometry and solely applying texture optimizations, this artifact can be alleviated (e.g., TexGen [22]).
Figure 16: Generated head avatars that can be manipulated (e.g., accessories) in graphics software without mesh post-processing.
Figure 17: More qualitative comparisons with recent methods that do not employ mesh representations for avatar generation.
Figure 18: Examples of texture transfer. Since the UV coordinates of the input template mesh are well preserved, various generated texture maps can be directly applied to the generated mesh.
Figure 19: Example of text-based texture editing. With extra text prompts, our method supports texture editing in partial regions.